The length of a rectangle is 3 more than twice its width. The perimeter is 48 feet. Find the width.
Answer by alicealc(293)

rectangle's length = L
rectangle's width = W

L = 3 + 2W
perimeter = 48
2L+ 2W = 48
2(3 + 2W) + 2W = 48
6 + 4W + 2W = 48
6W = 48 - 6
6W = 42
W = 42/6 = 7

so,the width of the rectangle is 7 feet